An aerospace company has submitted bids on two separate federal government defense contracts. The company president believes that there is a 41% probability of winning the first contract. If they win the first contract, the probability of winning the second is 75%. However, if they lose the first contract, the president thinks that the probability of winning the second contract decreases to 50%.
A. What is the probability that they win both contracts?
Probability =
B. What is the probability that they lose both contracts?
Probability =
C. What is the probability that they win only one contract?
Probability =

Answer by stanbon(75887)

An aerospace company has submitted bids on two separate federal government defense contracts. The company president believes that there is a 41% probability of winning the first contract. If they win the first contract, the probability of winning the second is 75%. However, if they lose the first contract, the president thinks that the probability of winning the second contract decreases to 50%.
A. What is the probability that they win both contracts?
Probability = P(W and W) = 0.41*0.75
B. What is the probability that they lose both contracts?
Probability = P(L and L) = 0.59*0.5
C. What is the probability that they win only one contract?
Probability = P(Win one) = 1 - [0.41*0.75+0.59*0.5]
=====================
Cheers,
Stan H.
--------------